OpenStudy (anonymous):

a simple pendulum is 50cm long its frequency of vibration on the surface of the earth is 1hz 0.5hz 0.7 0.65

OpenStudy (anonymous):

0.7 is answer....

OpenStudy (anonymous):

T=2*pi*squareroot(length/g) , T=2*3.14159*squareroot(50*10^-2/9.8)= 1.42s , f=1/T=1/1.42=0.7Hz